tva logoTennessee Valley Authority
.

Kingston Recovery Project

Dike C Weekly Progress Highlights

Each week, a status report is compiled to document the progress of time-critical work on dike C at the Kingston Recovery Project. Highlights of those reports are presented below along with a link to the week's complete dike C report in a PDF format.

Click here to view past updates from the dredging period, and past weekly recovery reports.

You may view a specific week’s report by selecting the date below:

Week Ending October 16, 2011

  • Air relief valves were installed on the diffusers. Work on all tasks associated with Dike C buttress construction was considered complete as of October 14, 2011.
  • Diffuser flow test is tentatively scheduled for the week of November 14.

View Complete Report (PDF)

Week Ending October 9, 2011

  • As-built drawings for the buttress are being prepared.
  • Crews cleaned and restored the peninsula and diffuser areas.
  • The flow test scheduled for October 10 is postponed until later in the year due to the plant outage and the pumps being offline.
  • Air relief valves still need to be installed on the diffusers.

View Complete Report (PDF)

Week Ending October 2, 2011

  • The diffusers were installed on Monday September 26. The temporary siphons were turned off at 6 PM on Monday.
  • The overflow pipe was installed on September 27.
  • A flow test on the newly installed diffusers is scheduled for the week of October 10.
  • Monitoring of the out flow from the Stilling Pond has been returned to the KIF Plant.

View Complete Report (PDF)

Week Ending September 18, 2011

  • The temporary siphons continue to function well.
  • Siphons will operate until diffuser installation is completed in late September. Two of the six
    diffusers have been installed.

View Complete Report (PDF)

Week Ending September 11, 2011

  • The temporary siphons continue to function well.
  • Siphons will operate until the diffusers are installed in mid- to late-September. Diffuser installation was delayed until the week of September 19 due to high water levels caused by recent excessive rains.

View Complete Report (PDF)

Week Ending August 28, 2011

  • TVA continued working on completing non-design related punch list items such as installing bollards around instruments.
  • The temporary siphons continue to function well.
  • Siphons will operate until the diffusers are installed in early September.

View Complete Report (PDF)

Week Ending August 7, 2011

  • Civil Projects (CP) completed placement of the rock buttress on Dike C.
  • The temporary siphons continue to function well.
  • Siphons will operate until the diffusers are installed in early September.
  • Began welding first sections of the new diffuser pipes.

View Complete Report (PDF)

Week Ending July 24, 2011

  • The temporary siphons continue to function well.
  • Civil Projects continued construction of the buttress in the intake channel.
  • Approximately 5,115 feet (96%) of the buttress has been completed.

Week Ending July 17, 2011

  • The temporary siphons continue to function well.
  • Completed removal of the causeway.
  • Construction of the buttress continued in the intake channel.
  • Approximately 5,022 feet (95%) of buttress have been completed.
  • View Complete Report (PDF)

Week Ending July 10, 2011

  • The temporary siphons continue to function well.
  • Civil Projects continued removing the causeway.
  • Regarding buttress construction, crews have finished the skimmer wall peninsula area and have moved to the intake channel area. Work is scheduled to be complete by September.

View Complete Report (PDF)

Week Ending July 3, 2011

  • The temporary siphons, which were monitored over the weekend, are working well.
  • Civil Projects continued removing the causeway using a crane and clamshell bucket.
  • Buttress construction remains approximately 93%complete. Crews are finishing up the Skimmer Wall area and moving to the Intake Channel area.

View Complete Report (PDF)

Week Ending June 19, 2011

  • The temporary siphons are working well. Crews monitored operations 24/7 with only minor adjustments needed.
  • Removal of the causeway continued. A crane and clamshell bucket are being used to remove the material. Turbidity generated by the excavation is being monitored and booms have been put in place to help contain material close to the work area.
  • Construction continued on the buttress in the Skimmer Wall Peninsula area; buttress construction is approximately 4,800 feet completed (91%). Work in the Skimmer Wall Peninsula area will progress slower than other areas because more work is required for each foot calculated as completed.

View Complete Report (PDF)

Week Ending June 12, 2011

  • The temporary siphons were put into operation and have been able to draw down the Stilling Pond and maintain the target elevation of 754 feet.
  • The existing diffuser pipes were removed and are currently being staged on the peninsula side of the Intake Channel.
  • Civil Projects began constructing the buttress in the Skimmer Wall peninsula area.
  • Buttress construction remains at 4,761 feet completed and is 90% complete.

View Complete Report (PDF)

Week Ending June 5, 2011

  • Work began to remove the causeway.
  • 4,761 feet of buttress have been completed. Work on the buttress is about 90% complete.
  • EPA approved the request for additional buttress along the Skimmer Wall.
  • Siphon installation continues with work at about 90% complete. The siphons should be operational early next week.
  • Bridge removal is complete and the subcontractor has demobilized.
  • EPA approved the Work Plan for Segment B Part 2.

View Complete Report (PDF)

Week Ending May 29, 2011

  • Civil Projects has completed installing buttress on the intake side of the Segment B.
  • 4,761 feet of buttress has been completed. The work is approximately 90% complete.
  • EPA is reviewing Stantec’s response to EPA’S request for additional buttress along the Skimmer Wall.
  • The EPA has approved the plan for Causeway Removal.
  • Siphon installation continues. The work is approximately 60% complete.
  • Bridge removal continues. The work is approximately 90% complete.
  • EPA is reviewing the Work Plan for Segment B Part 2.

View Complete Report (PDF)

Week Ending May 15 2011

  • Civil Projects is installing buttress on the intake side of Segment B.
  • The buttress is approximately 83% complete.
  • Stantec is preparing revised drawings for additional buttress along the Skimmer Wall.
  • The EPA is reviewing the plan for Siphon Construction.
  • Mobilization of materials for siphon construction began.
  • Three of four spans of the bridge have been removed.
  • RSI collected vibracore samples in the intake channel to identify impacts of potential change in current when the causeway is removed.

View Complete Report (PDF)

Week Ending May 8, 2011

View Complete Report (PDF)

Week Ending May 1, 2011

  • Civil Projects is constructing an access ramp to allow Segment B material to be placed from the toe up.
  • Segment A is completed to Station 120.
  • Approximately 4400 feet of 5280 is completed or about 83% complete.
  • EPA is requesting that the wall be extended further along the peninsula.
  • TVA is working on a plan for Siphon Construction which requires EPA approval.
  • Bridge removal continues.

View Complete Report (PDF)

Week Ending April 17, 2011

  • Completed approximately 4,400 of buttress. The buttress is approximately 83% complete.
  • EPA gave approval for the bridge demolition.
  • EPA gave approval for a portion of Segment B. Stantec provided a design package for the second part of Segment B. Review comments to be provided next week.
  • Continued with design for siphon installation and diffuser reconfiguration.

View Complete Report (PDF)

Week Ending April 10, 2011

  • Completed approximately 4,250’ of buttress. The buttress is approximately 80% complete.
  • EPA gave approval for the bridge demolition.
  • Provided response to EPA comments on Segment B.
  • Continued with the design for siphon installation and diffuser reconfiguration..

View Complete Report (PDF)

Document Updated December 5, 2013 3:56 PM

 

           
Content for id "future1" Goes Here
Content for id "future2" Goes Here
Content for id "future3" Goes Here